Maurice is building a 2 ft tall model of the Gateway Arch in St.
Louis, Missouri. If he is using a 3 in:78.75 ft scale, how tall is the actual arch?

Answer by TimothyLamb(4379)

---
scale ft/ft:
H/h = (78.75 ft)/(3 in * 1/12 ft/in)
H/2 = 78.75/(3/12)
H = 2 * 78.75 / (3/12)
---
H = 630 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php